### v1

# AmpleGCG: Learning a Universal and Transferable Generative Model of Adversarial Suffixes for Jailbreaking Both Open and Closed LLMs

Abstract
Warning: This paper contains potentially offensive and harmful text.
As large language models (LLMs) become increasingly prevalent and integrated into autonomous systems, ensuring their safety is imperative. Despite significant strides toward safety alignment, recent work GCG (Zou et al., 2023) proposes a discrete token optimization algorithm and selects the single suffix with the lowest loss to successfully jailbreak aligned LLMs. In this work, we first discuss the drawbacks of solely picking the suffix with the lowest loss during GCG optimization for jailbreaking and uncover the missed successful suffixes during the intermediate steps. Moreover, we utilize those successful suffixes as training data to learn a generative model, named AmpleGCG, which captures the distribution of adversarial suffixes given a harmful query and enables the rapid generation of hundreds of suffixes for any harmful queries in seconds. AmpleGCG achieves near 100% attack success rate (ASR) on two aligned LLMs (Llama-2-7B-chat and Vicuna-7B), surpassing two strongest attack baselines. More interestingly, AmpleGCG also transfers seamlessly to attack different models, including closed-source LLMs, achieving a 99% ASR on the latest GPT-3.5. To summarize, our work amplifies the impact of GCG by training a generative model of adversarial suffixes that is universal to any harmful queries and transferable from attacking open-source LLMs to closed-source LLMs. In addition, it can generate 200 adversarial suffixes for one harmful query in only 4 seconds, rendering it more challenging to defend. Our code is available at https://github.com/OSU-NLP-Group/AmpleGCG.
1 Introduction
With large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) demonstrating impressive performance on a wide range of tasks, more and more of them are being integrated into autonomous systems and deployed to solve real-world problems. Hence, improving their safety and trustworthiness has attracted significant attention (Yang et al., 2024; Tang et al., 2024; Mo et al., 2024) with a large amount of efforts (Xu et al., 2024; Yuan et al., 2024; Bhatt et al., 2023) focusing on developing safety guardrails.
Unfortunately, recently proposed jailbreaking attacks (Zeng et al., 2024; Shah et al., 2023; Shen et al., 2023; Yu et al., 2023; Zou et al., 2023) have been shown successful to circumvent the safety guardrails. For example, a pioneering method named GCG (Zou et al., 2023) proposes to automatically produce discrete adversarial tokens and demonstrates its effectiveness in jailbreaking various aligned LLMs. However, existing jailbreaking methods, including GCG, only generate one single adversarial prompt for one query and leave many vulnerabilities unexplored. Additionally, they tend to be time-consuming, e.g., GCG requires hours of optimization to curate one adversarial suffix (Jain et al., 2023).
In this work, we take GCG as the testbed due to their resilience to standard alignment and explore several intriguing questions: How can we find as many vulnerabilities of an LLM as possible? Is it feasible to build a generative model to directly produce many adversarial suffixes given any harmful queries within a short time? Answering these questions would significantly enhance our understanding of the LLM’s security posture (Section 2).
In addressing these questions, we first analyze why some suffixes, despite exhibiting relatively low loss during GCG optimization, might fail to jailbreak the target LLMs. With a pilot study on this, we discover that the loss associated with the first token is disproportionately high despite the overall low loss, indicating that the model is likely to pick a tone of refusal at the beginning of inference. It will lead to subsequent generation in a safe mode. This observation underscores that loss is not a reliable measure and only selecting the suffix with the lowest loss is not the best strategy. Based on this observation, we keep all candidate suffixes sampled during the GCG optimization and use them to jailbreak, which we refer to as augmented GCG. Results show that the attack success rate (ASR) can be substantially increased from $\sim$20to $\sim$80on the aligned Llama-2-7B-Chat and many more successful suffixes can be discovered for each query (Section 3).
Based on these findings, we proceed to pursue the two questions asked earlier. We propose to train a generative model, termed AmpleGCG, to model the mapping between a harmful query and its customized adversarial suffixes. We use a straightforward pipeline to collect training data from augmented GCG for developing AmpleGCG. To systematically evaluate the effectiveness of AmpleGCG, we carry out experiments to generate tailored suffixes for each harmful query in the test query set. Results reveal that our AmpleGCG could achieve near 100% ASR on both Vicuna-7B and Llama-2-7B-Chat by sampling around 200 suffixes, markedly outperforming the two strongest baselines as well as augmented GCG. This highlights AmpleGCG’s ability to uncover a wide range of vulnerabilities for different queries. Moreover, the suffixes generated by AmpleGCG exhibit greater diversity than those derived directly from augmented GCG with no overlap, pointing to the diverse and new coverage of vulnerabilities. Notably, AmpleGCG can produce 200 adversarial suffixes for each query in around 4 seconds only, indicating high efficiency (Section 4).
Even more interestingly, we find that AmpleGCG transfers well to different models. Specifically, AmpleGCG trained on open-source models exhibit remarkable ASRs on both unseen open-source and closed-source models. By simply adding specific affirmative prefixes (e.g., ”Sure, here is”) to the generated suffixes from AmpleGCG , ASR can reach 99% on the latest version of GPT-3.5, even higher than the ASR on the previous version (90%). This also issues a cautionary note regarding OpenAI’s update of more advanced models, highlighting that such advancements may come at the cost of compromised safety (Section 4).
One potential concern about GCG and AmpleGCG, where adversarial suffixes are unnatural, is that they can be easily detected by a perplexity-based defense mechanism. To address this concern, we show that by simply repeating a harmful query for multiple times at inference time, AmpleGCG’s generated adversarial suffixes can successfully evade perplexity-based defenses with an 80% ASR (Section 4).
To summarize, we learn a universal generative model that captures a distribution of customized adversarial suffixes given any harmful query and can generate many successful adversarial suffixes efficiently at inference time. It could also transfer to both unseen open-source and closed-source LLMs and bypass perplexity-based defense mechanisms.
2 Preliminaries
Background on GCG. GCG (Zou et al., 2023) is a pioneering approach to elicit objectionable content from aligned LLMs through discrete token-level optimization. Specifically, given a harmful query $x_{1:m}$, GCG aims to find a suffix $x_{m+1:m+l}$ and append it to the original query so as to form an adversarial query $x_{1:m+l}=concat(x_{1:m},x_{m+1:m+l})$. To make the victim model(s) prone to outputting harmful content given an adversarial query, GCG requires the response of the model(s) to start with positive affirmation $y$ (i.e., ”Sure, here is how to {$x_{1:m}$}”). Under this context, GCG leverages the standard autoregressive objective function as the loss function:
$\mathcal{L}(x_{m+1:m+l})=-\log p(y|concat(x_{1:m},x_{m+1:m+l}))=-\log p(y|x_{1:m+l})$ (1)
to optimize the adversarial suffix $x_{m+1:m+l}$ via proposed Greedy Coordinate Gradient (GCG) algorithm. Essentially, the algorithm randomly initializes the adversarial suffix and then samples a batch of candidate suffixes based on the gradient (Ebrahimi et al., 2018) at each optimization step. It uses the suffix with the lowest loss up to that point for the next iteration and continues the optimization until the number of max steps is reached. Ultimately, GCG only picks the suffix with the lowest loss throughout the entire optimization process.
Besides, two settings were designed by Zou et al. (2023), one for finding an adversarial suffix for each individual query and the other for finding a universal suffix for all queries: (1) The individual query setting intends to optimize the objective function for each harmful query and obtain one tailored adversarial suffix for it. (2) The mutliple queries setting is designed for optimizing multiple harmful queries simultaneously by fusing the gradients and generates a universal adversarial suffix for multiple queries together. Both settings are compatible to using one or more open-source victim models during optimization.
Under either setting, the effectiveness of an optimized suffix is measured by detecting the harmfulness of the output from victim models with input being the adversarial query $x_{1:m+l}$.
Motivation. GCG has achieved remarkable success in jailbreaking aligned LLMs and has been employed as an automated red teaming tool (Mazeika et al., 2024). Besides, GCG stands our among other jailbreaking strategies by targeting long-tail gibberish cases, which rarely occur in real-world scenarios and therefore cannot be trivially mitigated through conventional human alignment techniques (Ouyang et al., 2022), even with more human alignment data. It is also why we take GCG as our testbed for further exploration.
However, GCG settles on only one adversarial suffix with the lowest loss to attack the victim models, leading to the oversight of many adversarial suffixes and making them easy to fix due to the limited number. In this paper, we aim to explore several intriguing questions: (1) Is loss really a suitable metric to select a potential adversarial suffix for jailbreaking? (2) How can we find as many vulnerabilities as possible for each specific query? Is it possible to build a generative model to directly produce a vast array of customized suffixes given a harmful query within a short time, instead of going through GCG optimization which is time-consuming? Addressing these questions could help us better understand the safety posture of LLMs and further amplify the impact of GCG either as an approach for jailbreaking or a tool for automated red-teaming by unveiling more vulnerabilities of LLMs.
3 Rediscovering GCG: Loss Is Not a Good Reference for Suffix Selection
In this section, we take a deep look at GCG and find that for each suffix, its corresponding value of the loss function in GCG is not a good reference of its jailbreaking performance (success or failure). Based on this finding, we then propose augmented GCG, which collects all suffixes sampled at each optimization step and uses them to attack a victim model.
Benchmark. We primarily use AdvBench Harmful Behaviors introduced by (Zou et al., 2023), which comprises 520 instances. We adopt their harmful behavior setting in our experiments unless otherwise stated.
Hyperparameter Setup. We adhere to GCG’s standard optimization procedure, applying 500 steps for all victim models, with the exception of 1000 steps for Llama-2-7B-Chat. At each step, we sample 256 candidate suffixes. When optimizing a suffix, we keep its default length of 20 from GCG. For each victim model, we obtain the output by greedy decoding and set the maximum number of tokens at 100 because using 100 tokens could achieve a similar ASR with using even longer tokens (Mazeika et al., 2024).
Harmfulness Evaluator. Different from GCG, we employ two methods to jointly measure ASR. The first method is a string-based evaluator, which assesses harmfulness by checking the presence of predefined keywords in the responses of the victim models. The full list of keywords used for evaluation can be found in Appendix P. To more reliably classify the harmfulness of a response  (Qi et al., 2023), we further include the Beaver-Cost model (Dai et al., 2023) as a model-based evaluator which achieves 95% human agreement in the report. Generally, a response of a victim model is regarded as harmful if the above two evaluators detect harmful content in it unless otherwise stated. We calculate the ASR metric as the average success rate across all queries.
Figure 1: The $\log$(Loss) [^2] We apply the logarithmic operation to the loss to more clearly illustrate its decreasing trend.of candidate suffixes during GCG optimization over Llama-2-7B-Chat for one randomly sampled query. Red points/stars denote successful adversarial suffixes; blue points/stars indicate failed ones. The star at each step represents the anchor suffix with the lowest loss up to that point and is used for the next step of optimization.
3.1 Loss is Not a Good Reference for Suffix Selection
GCG uses the loss function to select the suffix with the lowest loss to jailbreak. Is loss really a good indicator of jailbreaking performance? To answer this question, we visualize the losses of candidate suffixes during the optimization steps, where we randomly select one harmful query from AdvBench and carry out the optimization process under the individual query setting using Llama-2-7B-Chat as the victim model. From Figure 2, we observe that (1) a suffix with a low loss might fail to jailbreak and (2) there are many successful suffixes with higher losses at each step, even though suffixes with a lower loss have a higher chance to jailbreak. We present the optimization visualization for Vicuna-7B in Appendix I, which shows more unexplored successful suffixes in 500 steps.
Why might a suffix corresponding to a low loss fail to jailbreak? This is related to the known issue of exposure bias in training autoregressive models in a teacher-forcing fashion (Bengio et al., 2015). For suffixes that correspond to a low loss but fail to jailbreak, despite the overall low loss across all tokens in the target output, we find that the loss associated with the first token is substantially higher compared to tokens at subsequent positions, indicating the optimization over the first token is not successful. When such optimized suffix is used to jailbreak the victim model, the model tends to pick a refusal tone for the first token (e.g., ”Sorry”). Since autoregressive models’ decoding at inference time depends on previously generated tokens, the subsequent tokens generated after the first refusal token tend to stay in a safe mode. Experiments of a controlled teacher-forcing generation in Appendix M includes more detailed discussions and analysis on this observation.
3.2 Augmenting GCG with Overgeneration
Taking previous empirical findings in mind, we keep all candidate suffixes (which we refer as overgeneration) sampled during the GCG optimization and present the effectiveness of GCG augmented with overgeneration, dubbed as augmented GCG (Algorithm 1).
Experimental Setting. We randomly pick 30 harmful queries from AdvBench to evaluate if augmented GCG could further improve ASR and find more successful suffixes. For each harmful query, as long as one suffix succeeds, the attack is deemed successful.
(1) For the individual query setting, we evaluate all possible overgenerated candidates during optimization (all points in Figure 2). (2) For the mutliple queries setting, we randomly pick 25 harmful queries (which we ensure are non-overlapping with the above 30 test queries) as our training set and optimize for a universal suffix. Note that both individual query and mutliple queries are compatible with the overgeneration strategy. To reduce the memory and time cost, in this mutliple queries setting, we only save the suffixes with the lowest loss (stars in Figure 2). GCG’s default setting, picking the suffix with the lowest loss at the entire optimization (purple triangle in Figure 2), will be our baseline to compare with.
Besides evaluating ASR, we also assess the number of Unique Successful Suffixes (USS) identified for each jailbroken query on average. While ASR intends to measure the success rate across queries, the USS number reveals the total number of the identified adversarial suffixes for each query on average. These metrics together are crucial for gaining a thorough understanding of an attack method’s effectiveness and the target LLM’s safety posture.
Experimental Setting Llama-2-7B-Chat Vicuna-7B
ASR(%) USS ASR(%) USS
Individual Query GCG default 20.00 1 93.33 1
Overgenerate 76.66 12320.86 100.00 46204.66
Multiple Queries GCG default 23.00 1 93.33 1
Overgenerate 83.33 129 100.00 83
Table 1: The ASR and USS results for Llama-2-7B-Chat and Vicuna-7B across 30 test queries demonstrate substantial improvements with overgeneration.
Results. Our results are shown in Table 1. For both individual query and mutliple queries settings, overgeneration boost ASR from $\sim$20to $\sim$80for Llama-2-7B-Chat and push it to 100% for Vicuna-7B. Besides, with a higher USS, it discovers more adversarial suffixes for each query compared to the default GCG setting. Due to memory and time constraints, we only pick the suffix with the lowest loss at each step under the mutliple queries setting (which is why USS is smaller in this setting). ASR and USS could be further enhanced if we extensively sample all candidates like the individual query setting. This demonstrates that many successful suffixes are overlooked by solely picking the suffix with the lowest loss and, more importantly, overgeneration amplifies the effectiveness of GCG by exposing more vulnerabilities of the target LLMs. With more discovered adversarial suffixes and vulnerabilities, augmented GCG could also help to conduct automated red-teaming more effectively than default GCG.
4 AmpleGCG: Learning a Universal and Transferable Generative Model
Now an intriguing question arises: How can we encompass a broader range of potentially successful suffixes to jailbreak the target model(s)? In this section, we propose training a universal generative model (dubbed as AmpleGCG) by collecting data from a simple overgenerate-then-filter pipeline and learning the distribution of adversarial suffixes for any harmful queries. With the learned AmpleGCG, we could rapidly sample hundreds of adversarial suffixes for a harmful query in seconds. we also demonstrate its efficacy in transferring across both open and closed LLMs, countering perplexity defense strategies.
4.1 Methodology
Overgenerate-then-Filter Pipeline. Given that there are a number of successful suffixes missed during the GCG optimization, we propose a straitforward overgenerate-then-filter (OTF) pipeline, extended on the overgeneration strategy (augmented GCG) described in Section 3. Specifically, we apply the model-based and string-based evaluators to filter out failed suffixes for corresponding victim models after overgeneration. Illustration of OTF is included in the right part of Figure 3.
Data Collection and Training for AmpleGCG. We initially draw 445 queries from AdvBench and apply the OTF pipeline to collect successful adversarial suffixes for each query under the individual query setting. Since Llama-2-7B-Chat built with stronger defense mechanisms remains refusing certain queries even with the overgeneration strategy, we only retain 318 queries that have successful suffixes to both Llama-2-7B-Chat and Vicuna-7B as our training queries. The remaining 127 (i.e., 445--318) queries, which fail in breaching Llama-2-7B-Chat with OTF, are identified as challenging. We deliberately incorporate a subset of these to create a challenging test set for subsequent evaluation.
With the collected training pairs (<harmful query, adversarial suffix>), we adapt Llama-2-7B (Touvron et al., 2023) as our base model for AmpleGCG and train it to generate adversarial suffixes given a harmful query. We ablate a number of decoding approaches during inference to generate suffixes from AmpleGCG and decide to take group beam search (Vijayakumar et al., 2016) as our default decoding method to encourage generating diverse suffixes at each sampling time. Specifically, we set the number of groups the same as the number of beams and set diversity penalty to 1.0333Parameters are defined in https://huggingface.co/docs/transformers/v4.18.0/en/main_classes/text_generation. The right part of Figure 3 illustrates how we use the collected data from the OTF pipeline to train AmpleGCG and more details can be found in Appendix B.
4.2 Experimental Setup
Test Set. We select 100 harmful queries from AdvBench and ensure no overlapping with the aforementioned 318 training queries. In particular, 56 of them are randomly picked from the challenging 127 cases that fail to jailbreak Llama-2-7B-Chat under augmented GCG’s individual query setting and 44 queries are randomly selected from queries that are not involved in creating training sets.
Baselines. We compare our AmpleGCG with the GCG default setting under both the individual query and mutliple queries settings and its augmented counterpart. According to (Mazeika et al., 2024), AutoDAN (Liu et al., 2023c) serves as the second most effective approach to jailbreaking after GCG, we report the performance of AutoDAN-GA as well.
Target Models. We present the effectiveness of different attack methods in jailbreaking Vicuna-7B (Chiang et al., 2023) and Llama-2-7B-Chat (Touvron et al., 2023), given any harmful test queries. To show AmpleGCG’s transferability, we further include open-source Mistral-7B-Instruct (Jiang et al., 2023) and closed-source GPT-3.5-0613, GPT-3.5-0125, GPT-4-0613 models (OpenAI et al., 2023) as our target models.
4.3 Results
Effectiveness, Efficiency, and Broad Coverage of AmpleGCG. Table 4.3 and Table H (in Appendix H) present the results of jailbreaking Llama-2-7B-Chat and Vicuna-7B, respectively. We observe that AmpleGCG could obtain a high ASR compared to the two most effective baselines, GCG default setting and AutoDAN, by a substantial margin. Even when GCG is augmented with overgeneration, obtaining 257 and 87 universal suffixes for Llama-2-7B-Chat and Vicuna-7B under the mutliple queries setting respectively, AmpleGCG attains a higher ASR with fewer samples, achieving up to 99% ASR.
Table 2: Results of different jailbreaking methods over Llama-2-7B-Chat on 100 test queries. The number in the bracket means the number of sampled suffixes from either AmpleGCG or augmented GCG.
Experimental Setting Llama-2-7B-Chat
Methods Sampling Strategies ASR(%) USS Diversity Time Cost
GCG individual query GCG default   (-) 1.00 1 - 122h
Overgenerate (192k) 16.00 12k 28.31 122h
[…]
Overgenerate (257) 94.00 30.12 50.74 6h
AutoDAN - 40.00 1 - 14m
AmpleGCG from Llama-2-7B-Chat Group Beam Search (50) 83.00 4.46 47.79 50s
Group Beam Search (100) 93.00 6.92 59.56 3m
Group Beam Search (200) 99.00 11.03 69.98 6m
Group Beam Search (257) 99.00 11.94 71.33 7m
Group Beam Search (1000) 100.00 37.66 78.58 35m
Additionally, AmpleGCG uncovers more vulnerabilities with higher USS and more diverse adversarial suffixes—where diversity is assessed by average pairwise edit distances between all generated nonsensical suffixes—compared to both the default GCG and augmented GCG under the mutliple queries setting. Though augmented GCG can obtain 192k suffixes with 122h computation under the individual query setting, their diversity is relatively low, and it can’t find vulnerabilities in a diverse span like AmpleGCG. We also automatically verify that the suffixes from AmpleGCG are distinct from augmented GCG’s 192k instances, highlighting the new coverage of vulnerabilities from AmpleGCG and complementing the vulnerabilities found by augmented GCG.
Notably, it only takes AmpleGCG 6 mins [^4] Time cost is derived by running on a single NVIDIA A100 80GB GPU with AMD EPYC 7742 64-Core Processor. in total to produce 200 suffixes for each of the 100 test queries (4 seconds per test query) that reach near 99% ASR, substantially reducing the time cost of curating adversarial prompts for LLMs (Jain et al., 2023).
Given that the test set comprises 56 challenging queries that fail to jailbreak Llama-2-7B-Chat using augmented GCG under individual query setting (Section 4.1), we particularly study if AmpleGCG, trained on successful queries, could jailbreak these 56 challenging failure queries to demonstrate its in-domain (IND) generalization.
Experimental Setting Challenging Queries
Methods Sampling Strategies ASR(%)
GCG individual query GCG default   (-) 0.00
Overgenerate (192k) 0.00
GCG mutliple queries GCG default   (-) 26.78
Overgenerate (257) 80.36
AmpleGCG from Llama-2-7B-Chat Group Beam Search (100) 96.42
Group Beam Search (1000) 100.00
Table 3: ASR results on challenging IND queries.
As shown in Table 3, AmpleGCG can jailbreak these challenging queries at 100% ASR that augmented GCG under individual query setting could not jailbreak at all. This is because AmpleGCG successfully learns the mapping between adversarial suffixes and queries, enabling it to generalize to challenging queries.
Moreover, we adapt MalicisousInstruct (Huang et al., 2023) (details in Appendix J) as an out-of-distribution (OOD) evaluation set to evaluate AmpleGCG’s OOD generalization. We only include GCG’s mutliple queries setting as our baseline here due to lower efficiency and efficacy of individual query setting as shown previously. Results are presented in 4.
Experimental Setting Victim Models
Llama-2-7B-Chat
Methods Sampling Strategies ASR(%)
GCG mutliple queries GCG default   (-) 39.00
Overgenerate (257) 80.00
AmpleGCG from Llama-2-7B-Chat Group beam search (100) 90.00
Group beam search (200) 99.00
Table 4: AmpleGCG shows generalization to OOD harmful categories and unseen formats from MaliciousInstruct.
AmpleGCG can outperform both the default GCG and augmented GCG using fewer samples of adversarial suffixes and reach 99% ASR, which indicates that AmpleGCG generalizes to different OOD harmful categories and different interrogative query formats from MaliciousInstruct.
In order to understand the effect of different decoding methods, we ablate three different decoding approaches in Appendix L. Our findings reveal that group beam search excels in exploring different adversarial suffixes compared to beam search and top_p decoding approaches, thus unlocking the full efficacy of AmpleGCG.
Transferability of AmpleGCG. We first investigate the transferability of AmpleGCG to attack open-source models that are not involved during the training process. Results (Appendix D) show that AmpleGCG could transfer well to different open-source models. Specifically, AmpleGCG trained based on the data from Llama-2-7B-Chat could achieve 72% transfer ASR on Vicuna-7B and Mistral-7B-Instruct averagely, surpassing semantic-keeping prompt from AutoDAN and augmented GCG. More analysis are included in Appendix D.
More excitingly, we delve into whether AmpleGCG, trained on open-source models, can be applied to attacking closed-source models. We collect training pairs from four models—Vicuna-7B, Vicuna-13B, Guanaco-7B, and Guanaco-13B (Dettmers et al., 2023), following the findings in (Zou et al., 2023). For details on our data collection methodology when optimizing across multiple models, please refer to Appendix B. For API cost consideration, we randomly select 50 queries from the 100 queries test set. We further incorporate two extra stringent evaluators for detecting harmfulness to reduce false positives: classifier from (Mazeika et al., 2024) and GPT-4 evaluator (evaluation prompt is in Appendix O). We take the GCG mutliple queries setting as our baseline, which is optimized using the aforementioned four models as well.
Methods Sampling Strategies + Tricks GPT-3.5-0613 GPT-3.5-0125 GPT-4-0613
ASR(%) ASR(%) ASR(%)
GCG mutliple queries GCG default   (-) 6.00 12.00 0.00
Overgenerate (140) 28.00 56.00 2.00
AutoDAN* (Llama-2-7B-Chat) - 2.00 0.00 0.00
AutoDAN* (Vicuna-7B) - 4.00 0.00 0.00
AmpleGCG Group beam search (200) 74.00 92.00 6.00
Group beam search (200) + AF 82.00 99.00 6.00
Group beam search (400) 80.00 98.00 12.00
Group beam search (400) + AF 90.00 99.00 8.00
Table 5: Transferability of different methods to attacking closed-source models.”AF” represents affirmative prefixes. * represents the source model that AutoDAN optimized on. Since AutoDAN is not compatible to a multiple-model optimization process, we follow the default AutoDAN implementation for two individual models and produced adversarial prompts are transferred to other victim models.
Compared to the default GCG, GCG augmented with overgeneration increases ASR by 4 times with 140 generated suffixes. By increasing the sampling times, AmpleGCG notably enhances the ASR performance, reaching up to 98% on the latest GPT-3.5-0125 model. To further boost this performance, we incorporate an affirmative prefix (”Sure, here is”)—denoted as AF in Table 5—at the end of the generated suffixes. This modification leads to an additional increase in ASR for both GPT-3.5-0613 and GPT-3.5-0125 models, with the latter nearly achieving a 100% ASR. This increase could be attributed to the models’ tendency to overlook their internal defense mechanisms after processing effective gibberish, and subsequently become unaligned models devoid of guardrails. Similar observations have been made by (Nasr et al., 2023), which noted that appending lengthy nonsensical tokens could bypass safeguards and retrieve private information from GPT-3.5. We include an example of attacking GPT-3.5-0125 successfully in Appendix N.To avoid false positives from model-based evaluators, we also manually verify the harmfulness of the generated content previously identified as harmful, according to the same principles used for GPT-4 evaluator. Given 99% ASR on the latest GPT-3.5 compared to 90% on the earlier version, we suspect that enhancing capabilities come at the expense of compromising safety unfortunately.
However, due to the stronger internal safeguard mechanisms of GPT-4, all approaches, including AmpleGCG, fail to achieve a decent ASR, while AmpleGCG’s performance still surpasses all baselines. We hypothesize that by increasing the sampling times, ASR could be further improved for AmpleGCG. Future work could also leverage more stringent harmfulness evaluators (such as GPT-4 or human evaluator) to collect training data for AmpleGCG to enhance its transferablity to attacking GPT-4.
AmpleGCG against Perplexity-based Defense. As suggested by (Jain et al., 2023), gibberish suffixes would be easily detected by perplexity-based detectors. To study the effectiveness of AmpleGCG against perplexity-based defenses, we propose two tricks to repeat the original query and extend the overall length, thereby leading to lower perplexity: (1) AmpleGCG Input Repeat (AIR) generates the corresponding suffixes from repetitive queries, i.e., $AmpleGCG(repeat(x_{1:m},j))$, where $repeat(x,j)$ means repeat the original query $x$ for $j$ times. Then, it appends the suffix to the repetitive queries to jailbreak while maintaining low perplexity. (2) We notice that AmpleGCG is only trained on non-repetitive queries and may lead to distribution shift if the input is repetitive. We further propose AmpleGCG Input Direct (AID) to directly generate suffixes via $AmpleGCG(x_{1:m})$, similarly to training procedure, but append this suffix to the repetitive queries. Results in Appendix F indicate that AmpleGCG successfully evades the defense with an 80% ASR using only 100 suffixes, outperforming GCG and AutoDAN, which achieve ASRs of 0% and 42%, respectively. Essentially, AmpleGCG extends its effectiveness to unfamiliar repetitive patterns, suggesting that the suffixes AmpleGCG produces are effective across different formats of queries. We also hypothesize that proficiency can be attributed to the recency bias issue (Liu et al., 2023a) of LLMs, only focusing on the last several statements while ignoring the previous part.
Table 2: Results of different jailbreaking methods over Llama-2-7B-Chat on 100 test queries. The number in the bracket means the number of sampled suffixes from either AmpleGCG or augmented GCG.